16. Solve the IVP with a Laplace transform method x + 2x + 2x = e-t, x(0) = 1, x(0) = 1
Answer #1

Given à + 22+2x =e-t Taking L T. get both hides or t?L??)= +2L?xS= (428+2)X(4) (s+3)(s+!)+H (r42s+2) (str) (:xl0i=1)l0)=,ATB = 1 2A+B+D =4 2 A+D - 4 Solwing we get A-|,B=0,D= in 3 %3D Putting thene then taking get valner inrerne omo L.7. ze57=L3